In Nyeri town there are only two milk processors. The local inverse demand for milk is given by: Q = 120− P, where P denotes price, Q denotes the total quantity measured in cartons. Both milk processors have the same cost function given by C = 30Q, where C is total cost and Q is output measured in cartons. What is the industry output (Q)?


1
Expert's answer
2022-01-18T10:09:21-0500

Given,

"Q=120 - P"

"P=120 - (Q_1+Q_2)"

"C1 = 30(Q_1) , C2 = 30(Q_2)"

for Firm 1,

"\pi_1=TR_1-TC_1=[{120-(Q_1+Q_2)}]Q_1-30Q_1"

"=(120)Q_1-(Q_1)^2-(Q_2)Q_1-30Q_1"

"=9Q_1-(Q_1)^2-(Q_2)Q_1"

"\frac{\delta\pi_1}{\delta Q_1}=90-2Q_1-Q_2=0"

"90=2Q_1+Q_2" ................eq.1

For Firm 2,

"\pi_2=TR_2-TC_2=[{120-(Q_1+Q_2)}]Q_2-30Q_2"

"=(120)Q_2-(Q_2)^2-(Q_1)Q_2-30Q_2"

"=9Q_2-(Q_2)^2-(Q_1)Q_2"

"\frac{\delta\pi_2}{\delta Q_2}=90-2Q_2-Q_1=0"

"90=2Q_2+Q_1" ................eq.2


On Solving eq.1 and eq.2,

we get ;-

"Q_2=30 , Q_1=30"

hence, each of them produce 30 units.